# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 5.1 SQ Lower Bounds for Learning Multi-Index Models
Assume the existence of an SQ algorithm $\mathcal{A}$ that given access to a distribution $D^{\prime}$ over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, returns a hypothesis $h$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[h(\mathbf{x})\neq y]<\tau-2\mathrm{OPT}$, with $\mathrm{OPT}=\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[f(\mathbf{x})\neq y]$. Using $\mathcal{A}$, we can reduce the testing problem to the learning problem.
Note that by running algorithm $\mathcal{A}$ to distribution $D_{\mathbf{P}}\in\mathcal{D}$, $\mathcal{A}$ returns a hypothesis $h$ with error $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{\mathbf{P}}}[h(\mathbf{x})\neq y]<\tau-2\mathrm{OPT}$. However, running a $\mathcal{A}$ with SQ oracle access to $D_{0}$ would result in a hypothesis $h$ that cannot perform better than $\tau-\mathrm{OPT}$, as $D_{0}$ is a product distribution for each fixed projection of $\mathbf{x}$ in $V$. As $\tau-\mathrm{OPT}>\tau-2\mathrm{OPT}$, we have that with an additional query of accuracy less than $\mathrm{OPT}$, we can solve $\mathcal{B}(\mathcal{D},D_{0})$. ∎
@@ 5.2 SQ Lower Bounds for Learning Multiclass Linear Classifiers with RCN
In this section, we present SQ hardness results for learning Multiclass Linear Classifiers in the presence of RCN under the Gaussian distribution. In particular, we show that achieving $\mathrm{OPT}+\epsilon$ error in the SQ model requires, either a super-polynomial number of queries, or queries with accuracy that depends on the spectral properties of the confusion matrix (see Definition 4.8), rather than on the amount of noise. The main result of this section is Theorem 5.3.
Theorem 5.3 (SQ Lower Bound for Learning under RCN).
@@
To construct our family of hard instances, we first define a two-dimensional model $f:\mathbb{R}^{2}\to[K]$ as follows:
@@
Geometrically, the classifier $f$ partitions the plane into $K$ equal angular sectors, assigning each sector to a different class. Now, given $\mathcal{Q}$, a set of $2\times d$ matrices, we can define a class of functions from $\mathbb{R}^{d}$ to $[K]$ as
@@
Before proving Theorem 5.3, we first need to show that we can match the appropriate number of moments with the standard Gaussian.
@@
    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}\left[y=i\mid\mathbf{x}\right]-{1}/{K}\right)p(\mathbf{x})\right]=0$ for all $i\in[K]$ ,
where $D$ is the distribution over $\mathbb{R}^{2}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ that satisfies Definition 4.8 for the confusion matrix $\mathbf{H}$ and the function $f$ as defined in Equation 6.
@@
Let $\mathbf{H}\in\mathbb{R}^{K\times K}$ be a doubly stochastic matrix and define the functions $\tilde{f}_{i}$ for $i\in[K]$ as $\tilde{f}_{i}(\mathbf{x})=\sum_{j=1}^{K}\mathbf{H}_{j,i}\mathds{1}(f(\mathbf{x})=j)$. For any function $p:\mathbb{R}^{2}\to\mathbb{R}$, we denote by $\bar{p}$ the zero mean centering of $p$, i.e., $\bar{p}(\mathbf{x})=p(\mathbf{x})-\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[p(\mathbf{x})\right]$, and by $\mathbf{v}^{(p)}\in\mathbb{R}^{K}$ the vector defined as $\mathbf{v}^{(p)}_{i}=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\bar{p}(\mathbf{x})\mathds{1}(f(\mathbf{x})=i)\right]$. We will refer to $\mathbf{v}^{{(p)}}$ as the vector of expectations of $p$. Also denote by $\boldsymbol{\omega}^{(k)}$ the root of unity vector of rate $k$, i.e. ${\boldsymbol{\omega}}^{(k)}\in\mathbb{C}^{K}$ such that $\boldsymbol{\omega}_{j}^{(k)}=e^{\frac{2\pi kji}{K}}$.
First, we show that a function $p$ has zero correlation with $\tilde{f}_{i}-1/K$ for all $i\in[K]$ under the standard Gaussian if and only if the vector of expectations of $p$ over the level sets belongs in the kernel of $\mathbf{H}^{\top}$.
@@
Let $p:\mathbb{R}^{2}\to\mathbb{R}$, then we have that $\mathbf{v}^{(p)}\in\ker(\mathbf{H}^{\top})\text{ if and only if }\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}[(\tilde{f}_{i}(\mathbf{x})-1/K)p(\mathbf{x})]=0\text{ for all $i\in[K]$}\;.$
@@
First note that from the rotational invariance of the standard Gaussian distribution we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}^{2}}[f(\mathbf{x})=i]=1/K$, for all $i\in[K]$. Hence, we have that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}[\tilde{f}_{i}(\mathbf{x})]=\sum_{j=1}^{K}\mathbf{H}_{ji}\operatorname*{\mathbf{E}}[\mathds{1}(f(\mathbf{x})=i)]=1/K$, from the assumption that $\mathbf{H}$ is column stochastic. Thus, for all $i\in[K]$ we have that
@@
Therefore, substituting the definition of $\tilde{f}_{i},i\in[K],$ we have that
@@
We next note that that all polynomials of degree at most $m$ can be expressed as a linear combination of polynomials of the form $(x+iy)^{a}(x-iy)^{b}$, where $a,b>0,a+b\leq m$. In the following claim, we show that the vectors associated with each polynomial of this basis are precisely the root of unity vectors.
@@
For $p(x,y)=(x+iy)^{a}(x-iy)^{b},a,b\in\mathbb{Z}_{+}$. If $a\neq b$, then there exists a constant $c\in\mathbb{C}$ such that $\mathbf{v}^{(p)}=c\boldsymbol{\omega}^{(a-b)}$; and if $a=b$, then $\mathbf{v}=\mathbf{0}$.
@@
First, note that by the definition of $f$ the region $\mathds{1}(f(x,y)=j)$ is equivalent to $\theta\in\left[{2\pi j}/{K}-{\pi}/{K},{2\pi j}/{K}+{\pi}/{K}\right]$, where $\theta=\arctan\left({y}/{x}\right)$. Fix $a,b\in\mathbb{Z}_{+},a\neq b$, and let $\theta_{1,j}={2\pi j}/{K}-{\pi}/{K}$, $\theta_{2,j}={2\pi j}/{K}+{\pi}/{K}$, for $j\in[K]$, and $p(x,y)=(x+iy)^{a}(x+iy)^{b}$. We have that
@@
Therefore, note that $\mathbf{v}_{(j+1)\mod K}^{(p)}=\mathbf{v}_{j}^{(p)}e^{\frac{2\pi i}{K}(a-b)}$ for all $j\in\{0,\dots,K-1\}$, since $\theta_{1,(j+1)}=\theta_{1,j}+{2\pi}/{K}$ and $\theta_{2,(j+1)}=\theta_{2,j}+{2\pi}/{K}$. Moreover, if $a\neq b$, we observe that the integral of $p(x,y)$ over the entire region is zero, which concludes the proof. ∎
Hence, by Claim 5.6, in order to have 00 correlation with all polynomials of degree at most $m$, we need to have that $\boldsymbol{\omega}^{(k)}\in\ker(\mathbf{H}^{\top})$ for all $k\in\{1,\dots,m,K-m,\dots,K-1\}$, as the roots of unity are symmetric, i.e., $\boldsymbol{\omega}^{(-k)}=\boldsymbol{\omega}^{(K-k)}$.
We first construct a stochastic vector $\mathbf{h}\in\mathbb{R}^{K}$ with $\mathbf{h}_{0}-\mathbf{h}_{j}=\Omega(1/K^{3})$, for all $j\in\{1,\dots,K-1\}$ that is also orthogonal to the aforementioned root of unity vectors for $m=\Omega(K)$. Next, we will extend this vector to a matrix. Define the vectors $\mathbf{h}^{(1)}=\boldsymbol{\omega}^{(0)}/K$, $\mathbf{h}^{(2)}=(\boldsymbol{\omega}^{(K/2-1)}+\boldsymbol{\omega}^{(K/2+1)})/K$ and $\mathbf{h}^{\prime}=2\mathbf{h}^{(1)}+\mathbf{h}^{(2)}$. Note that $\mathbf{h}^{(1)}_{j}=1/K$ and $\mathbf{h}_{j}^{(2)}=({2}/{K})(-1)^{j}\cos\left({2\pi j}/{K}\right)$ for all $j\in\{0,\dots,K-1\}$. Hence, we have that all the coordinates of $\mathbf{h}^{\prime}$ are nonnegative. Moreover, $\mathbf{h}^{\prime}$ attains its maximum value at $\mathbf{h}_{0}^{\prime}={4}/{K}$, since $\mathbf{h}_{K/2}^{\prime}=0$, because $(-1)^{j}\cos\left({2\pi j}/{K}\right)=-1$ when $j=K/2$ and $K$ is divisible by $4$. Furthermore, it holds that $\max_{j\in[K-1]}\mathbf{h}_{j}^{\prime}\leq 2/K+\max_{j\in[K-1]}\lvert\mathbf{h}_{j}^{(2)}\rvert=2/K+2/K\cos\left({2\pi}/{K}\right)$. As a result
@@
where the second inequality follows from the Taylor series $\cos(x)=1-\frac{x^{2}}{2!}+\frac{x^{4}}{4!}-\frac{x^{6}}{6!}+\cdots$, which is an alternating decreasing series when $x={2\pi}/{K}\leq 1$, thus it can be upper bounded by the first three terms. Let $\mathbf{h}$ be the normalization of $\mathbf{h}^{\prime}$ such that the sum of its coordinates equals $1$. Given that each element of $\mathbf{h}^{\prime}$ is at most $4/K$, we have that the normalization factor is at most $4$. Hence, $\mathbf{h}_{0}-\max_{j\in[K-1]}\mathbf{h}_{j}\gtrsim{1}/{K^{3}}$. Moreover, from the well-known fact that $\boldsymbol{\omega}^{(0)},\dots,\boldsymbol{\omega}^{(K-1)}$ form an orthogonal basis of $\mathbb{C}^{K}$, we have that $\mathbf{h}$ is orthogonal to $\boldsymbol{\omega}^{(k)}$ for all $k\in\{1,\dots,K/2-2,K/2,K/2+2,\dots,K-1\}$, since it is a linear combination of $\boldsymbol{\omega}^{(0)},\boldsymbol{\omega}^{(K/2-1)},\boldsymbol{\omega}^{(K/2+1)}$.
Now consider $\mathbf{H}$, the circulant matrix generated by row shifts of the vector $h$. By the definition of $\mathbf{H}$, we have that $\mathbf{H}$ is a row stochastic matrix and that $\min_{i\neq j}\mathbf{H}_{ii}-\mathbf{H}_{i,j}=\Omega(1/K^{3})$. Moreover, $\mathbf{H}$ is symmetric, because it holds that $\mathbf{H}_{i,j}=\mathbf{h}_{j-i\mod K}=\mathbf{h}_{i-j\mod K}=\mathbf{H}_{j,i}$ and $\mathbf{h}_{j}=\mathbf{h}_{K-j},j\in[K-1]$. This equality is justified by the identity $(-1)^{j}\cos\left({2\pi j}/{K}\right)=(-1)^{K-j}\cos\left({2\pi(K-j)}/{K}\right)$, which holds since $K$ is an even integer. Hence $\mathbf{H}$ is also column stochastic.
Finally, by A.6 we have that the eigenvectors of any circulant matrix are exactly $\boldsymbol{\omega}^{(k)}$ with eigenvalues $\mathbf{h}\cdot\boldsymbol{\omega}^{(k)},k\in\{0,\dots,K-1\}$. Hence, $\boldsymbol{\omega}^{(k)}\in\ker(\mathbf{H})=\ker(\mathbf{H}^{\top})$ for all $k\in\{1,\dots,K/2-2,K/2,K/2+2,\dots,K-1\}$. Therefore, by applying Claim 5.5, we conclude the proof of Lemma 5.4. ∎
@@
Consider the case where $K$ is divisible by $4$ and let $\mathbf{H}$ be a matrix that satisfies the statement of Lemma 5.4 Let $f:\mathbb{R}^{2}\to[K]$ be the linear classifier defined in Equation 6, and let $\mathcal{F}_{\mathcal{Q}}$ be the class of function defined in Equation 7, where $\mathcal{Q}\subseteq\mathbb{R}^{d\times 2}$ the family that satisfies the statement of C.4. We define the family of distributions $\mathcal{D}$ such that $D_{\mathbf{Q}}\in\mathcal{D}$ if and only if $D_{\mathbf{Q}}$ is a joint distribution $(\mathbf{x},y_{\mathbf{Q}})$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $y_{\mathbf{Q}}$ is a random variable supported on $[K]$ defined as $\operatorname*{\mathbf{Pr}}[y_{\mathbf{Q}}=i\mid\mathbf{x}]=\mathbf{H}_{f(\mathbf{Q}\mathbf{x}),i}$ for some $\mathbf{Q}\in\mathbf{\mathcal{missing}}{Q}$. Note that all distributions in $\mathcal{D}$ satisfy Definition 4.8 for confusion matrix $\mathbf{H}$ and some function in $\mathcal{F}_{\mathcal{Q}}$.

### v2

@@ 5.1 SQ Lower Bounds for Learning Multi-Index Models
Assume there exists an SQ algorithm $\mathcal{A}$ that, when given SQ access to any distribution $D^{\prime}$ over $\mathbb{R}^{d}\times\mathcal{Y}$ with standard Gaussian $\mathbf{x}$-marginals, outputs a hypothesis $h$ satisfying $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[h(\mathbf{x})\neq y]<\tau-2\mathrm{OPT}$, where $\mathrm{OPT}=\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[f(\mathbf{x})\neq y]$. Using such an algorithm $\mathcal{A}$, we show that we can solve the testing problem.
Suppose we run $\mathcal{A}$ on a distribution $D_{\mathbf{P}}\in\mathcal{D}$. Then, by definition, $\mathcal{A}$ returns a hypothesis $h$ with error $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{\mathbf{P}}}[h(\mathbf{x})\neq y]<\tau-2\mathrm{OPT}$. In contrast, when we run $\mathcal{A}$ with SQ oracle access to $D_{0}$ no hypothesis can achieve error lower than $\tau-\mathrm{OPT}$, as $D_{0}$ is a product distribution for each fixed projection of $\mathbf{x}$ in $V$. Since $\tau-\mathrm{OPT}>\tau-2\mathrm{OPT}$, an additional query of tolerance less than $\mathrm{OPT}$, would distinguish between $D_{\mathbf{P}}$ and $D_{0}$, thus solving $\mathcal{B}(\mathcal{D},D_{0})$. ∎
@@ 5.2 SQ Lower Bounds for Learning Multiclass Linear Classifiers with RCN
In this section, we present an SQ hardness result for learning Multiclass Linear Classifiers in the presence of RCN under the Gaussian distribution. In particular, we show that achieving $0-1$ error of $\mathrm{OPT}+\epsilon$ in the SQ model must either make a super-polynomial (with respect the dimension $d$) number of queries or queries with accuracy that depends on the spectral properties of the confusion matrix (see Definition 4.8), rather than on the amount of noise. The main result of this section is Theorem 1.9.
Theorem 5.4 (SQ Lower Bound for Learning under MLC RCN).
@@
To construct our family of hard instances, we first define a two-dimensional classifier $f:\mathbb{R}^{2}\to[K]$ as follows:
@@
Geometrically, the classifier $f$ partitions the plane into $K$ equal angular sectors, assigning each sector to a different class. Given a set of $2\times d$ matrices $\mathcal{Q}$, we can define a class of functions from $\mathbb{R}^{d}$ to $[K]$ as follows
@@
We show that it is hard to distinguish a distribution labeled by a classifier in $\mathcal{F}_{\mathcal{Q}}$ corrupted with RCN from a distribution with random labels independent of $\mathbf{x}$. Before proving Theorem 5.4, we first need to show that we can match the appropriate number of moments with the standard Gaussian.
@@
Let $K\in\mathbb{Z}_{+}$ be such that $K\geq 8$ and $K$ is divisible by $4$. Then there exists a $K\times K$ doubly stochastic matrix $\mathbf{H}$ and a distribution $D$ over $\mathbb{R}^{2}\times[K]$, whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, satisfying Definition 4.8 for the confusion matrix $\mathbf{H}$ and the function $f$ defined in Equation 6, such that:
@@
    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}\left[y=i\mid\mathbf{x}\right]-{1}/{K}\right)p(\mathbf{x})\right]=0$ for all $i\in[K]$
@@
Let $\mathbf{H}\in\mathbb{R}^{K\times K}$ be a doubly stochastic matrix and define the functions $\tilde{f}_{i}$ for $i\in[K]$ as $\tilde{f}_{i}(\mathbf{x})=\sum_{j=1}^{K}\mathbf{H}_{j,i}\mathds{1}(f(\mathbf{x})=j)$. For any function $p:\mathbb{R}^{2}\to\mathbb{R}$, we denote by $\bar{p}$ the zero mean centering of $p$, i.e., $\bar{p}(\mathbf{x})=p(\mathbf{x})-\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[p(\mathbf{x})\right]$, and by $\mathbf{v}^{(p)}\in\mathbb{R}^{K}$ the vector defined as $\mathbf{v}^{(p)}_{i}=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\bar{p}(\mathbf{x})\mathds{1}(f(\mathbf{x})=i)\right]$. We will refer to $\mathbf{v}^{{(p)}}$ as the vector of expectations of $p$. Also denote by $\boldsymbol{\omega}^{(k)}$ the root of unity vector of rate $k$, i.e. ${\boldsymbol{\omega}}^{(k)}\in\mathbb{C}^{K}$ such that $\boldsymbol{\omega}_{j}^{(k)}=e^{\frac{2\pi kji}{K}}$, where $i$ is the imaginary unit.
We start by showing that a function $p$ has zero correlation with $\tilde{f}_{i}-1/K$ for all $i\in[K]$ under the standard Gaussian if and only if the vector of expectations of $p$ over the level sets belongs in the kernel of $\mathbf{H}^{\top}$.
@@
Let $p:\mathbb{R}^{2}\to\mathbb{R}$. We have that $\mathbf{v}^{(p)}\in\ker(\mathbf{H}^{\top})$ if and only if for all $i\in[K]$ it holds $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}[(\tilde{f}_{i}(\mathbf{x})-1/K)p(\mathbf{x})]=0$.
@@
By the rotational invariance of the standard Gaussian we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}^{2}}[f(\mathbf{x})=i]=1/K$, for all $i\in[K]$. From the assumption that $\mathbf{H}$ is column stochastic, we have that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}[\tilde{f}_{i}(\mathbf{x})]=\sum_{j=1}^{K}\mathbf{H}_{ji}\operatorname*{\mathbf{E}}[\mathds{1}(f(\mathbf{x})=i)]=1/K$. Thus, for all $i\in[K]$ we have that
@@
Therefore, substituting the definition of $\tilde{f}_{i}$ we have that
@@
We next observe that all polynomials $\mathbb{R}^{2}\mapsto\mathbb{R}$ of degree at most $m$ can be expressed as a linear combination of polynomials of the form $(x+iy)^{a}(x-iy)^{b}$, where $a,b>0,a+b\leq m$ and $i$ is the imaginary unit. Specifically, we show that the vectors associated with each polynomial of this basis are precisely the root of unity vectors.
@@
Let $p(x,y)=(x+iy)^{a}(x-iy)^{b},a,b\in\mathbb{Z}_{+}$. If $a\neq b$, then there exists a constant $c\in\mathbb{C}$ such that $\mathbf{v}^{(p)}=c\boldsymbol{\omega}^{(a-b)}$; and if $a=b$, then $\mathbf{v}=\mathbf{0}$.
@@
By the definition of $f$, the region $\{(x,y)\in\mathbb{R}^{2}:f(x,y)=j\}$ is equivalent to $\theta\in\left[{2\pi j}/{K}-{\pi}/{K},{2\pi j}/{K}+{\pi}/{K}\right]$, where $\theta=\arctan\left({y}/{x}\right)$. Fix $a,b\in\mathbb{Z}_{+},a\neq b$, and let $\theta_{1,j}={2\pi j}/{K}-{\pi}/{K}$, $\theta_{2,j}={2\pi j}/{K}+{\pi}/{K}$, $j\in[K]$. We have that
@@
Therefore, it holds that $\mathbf{v}_{(j+1)\mod K}^{(p)}=\mathbf{v}_{j}^{(p)}e^{\frac{2\pi i}{K}(a-b)}$ for all $j\in\{0,\dots,K-1\}$, since $\theta_{1,(j+1)}=\theta_{1,j}+{2\pi}/{K}$ and $\theta_{2,(j+1)}=\theta_{2,j}+{2\pi}/{K}$. Moreover, if $a\neq b$, we observe that the integral of $p(x,y)$ over $\mathbb{R}^{2}$ is zero, which concludes the proof of the claim. ∎
By Claim 5.7, in order to have 00 correlation with all polynomials of degree at most $m$, we need to have that $\boldsymbol{\omega}^{(k)}\in\ker(\mathbf{H}^{\top})$ for all $k\in\{1,\dots,m,K-m,\dots,K-1\}$, as the roots of unity are symmetric, i.e., $\boldsymbol{\omega}^{(-k)}=\boldsymbol{\omega}^{(K-k)}$.
We first construct a stochastic vector $\mathbf{h}\in\mathbb{R}^{K}$ with $\mathbf{h}_{0}-\mathbf{h}_{j}=\Omega(1/K^{3})$, for all $j\in\{1,\dots,K-1\}$ that is also orthogonal to the aforementioned root of unity vectors for $m=\Omega(K)$. Next, we will extend this vector to a matrix. Define the vectors $\mathbf{h}^{(1)}=\boldsymbol{\omega}^{(0)}/K$, $\mathbf{h}^{(2)}=(\boldsymbol{\omega}^{(K/2-1)}+\boldsymbol{\omega}^{(K/2+1)})/K$ and $\mathbf{h}^{\prime}=2\mathbf{h}^{(1)}+\mathbf{h}^{(2)}$. Note that $\mathbf{h}^{(1)}_{j}=1/K$ and $\mathbf{h}_{j}^{(2)}=({2}/{K})(-1)^{j}\cos\left({2\pi j}/{K}\right)$ for all $j\in\{0,\dots,K-1\}$. Hence, we have that all the coordinates of $\mathbf{h}^{\prime}$ are nonnegative. Moreover, $\mathbf{h}^{\prime}$ attains its maximum value at $\mathbf{h}_{0}^{\prime}={4}/{K}$, since $\mathbf{h}_{K/2}^{\prime}=0$ (because $(-1)^{j}\cos\left({2\pi j}/{K}\right)=-1$ when $j=K/2$ and $K$ is divisible by $4$). Furthermore, it holds that $\max_{j\in[K-1]}\mathbf{h}_{j}^{\prime}\leq 2/K+\max_{j\in[K-1]}\lvert\mathbf{h}_{j}^{(2)}\rvert=2/K+2/K\cos\left({2\pi}/{K}\right)$, and hence we have
@@
where the second inequality follows from the Taylor series of $\cos(x)=1-\frac{x^{2}}{2!}+\frac{x^{4}}{4!}-\frac{x^{6}}{6!}+\cdots$, which is an alternating decreasing series when $x={2\pi}/{K}\leq 1$ (thus it can be upper bounded by the first three terms). Let $\mathbf{h}$ be the normalization of $\mathbf{h}^{\prime}$ such that the sum of its coordinates equals $1$. Given that each element of $\mathbf{h}^{\prime}$ is at most $4/K$, we have that the normalization factor is at most $4$. Hence, $\mathbf{h}_{0}-\max_{j\in[K-1]}\mathbf{h}_{j}\gtrsim{1}/{K^{3}}$. Moreover, from the well-known fact that $\boldsymbol{\omega}^{(0)},\dots,\boldsymbol{\omega}^{(K-1)}$ form an orthogonal basis of $\mathbb{C}^{K}$, we have that $\mathbf{h}$ is orthogonal to $\boldsymbol{\omega}^{(k)}$ for all $k\in\{1,\dots,K/2-2,K/2,K/2+2,\dots,K-1\}$, since it is a linear combination of $\boldsymbol{\omega}^{(0)},\boldsymbol{\omega}^{(K/2-1)},\boldsymbol{\omega}^{(K/2+1)}$.
Now consider $\mathbf{H}$, the circulant matrix generated by row shifts of the vector $h$. By definition, we have that $\mathbf{H}$ is a row stochastic matrix and that $\min_{i\neq j}\mathbf{H}_{ii}-\mathbf{H}_{i,j}=\Omega(1/K^{3})$. Moreover, $\mathbf{H}$ is symmetric, because it holds that $\mathbf{H}_{i,j}=\mathbf{h}_{j-i\mod K}=\mathbf{h}_{i-j\mod K}=\mathbf{H}_{j,i}$ and $\mathbf{h}_{j}=\mathbf{h}_{K-j},j\in[K-1]$. This equality is justified by the identity $(-1)^{j}\cos\left({2\pi j}/{K}\right)=(-1)^{K-j}\cos\left({2\pi(K-j)}/{K}\right)$, which holds since $K$ is an even integer. Hence $\mathbf{H}$ is also column stochastic.
By A.6, the eigenvectors of any circulant matrix are exactly $\boldsymbol{\omega}^{(k)}$ with eigenvalues $\mathbf{h}\cdot\boldsymbol{\omega}^{(k)},k\in\{0,\dots,K-1\}$. Hence, $\boldsymbol{\omega}^{(k)}\in\ker(\mathbf{H})=\ker(\mathbf{H}^{\top})$ for all $k\in\{1,\dots,K/2-2,K/2,K/2+2,\dots,K-1\}$. Therefore, by applying Claim 5.6, we conclude the proof of Lemma 5.5. ∎
@@
Consider the case where $K$ is divisible by $4$ and let $\mathbf{H}$ be a matrix that satisfies the statement of Lemma 5.5 Let $f:\mathbb{R}^{2}\to[K]$ be the linear classifier defined in Equation 6, and let $\mathcal{F}_{\mathcal{Q}}$ be the class of functions defined in Equation 7, where $\mathcal{Q}\subseteq\mathbb{R}^{d\times 2}$ is the family that satisfies the statement of C.4. We define the family of distributions $\mathcal{D}$ such that $D_{\mathbf{Q}}\in\mathcal{D}$ if and only if $D_{\mathbf{Q}}$ is a joint distribution $(\mathbf{x},y_{\mathbf{Q}})$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $y_{\mathbf{Q}}$ is a random variable supported on $[K]$ defined as $\operatorname*{\mathbf{Pr}}[y_{\mathbf{Q}}=i\mid\mathbf{x}]=\mathbf{H}_{f(\mathbf{Q}\mathbf{x}),i}$ for some $\mathbf{Q}\in\mathbf{\mathcal{missing}}{Q}$. Note that all distributions in $\mathcal{D}$ satisfy Definition 4.8 for confusion matrix $\mathbf{H}$ and some function in $\mathcal{F}_{\mathcal{Q}}$. Denote by $D$ the product distribution with $\mathbf{x}$-marginals distributed as standard normal and the labels $y$ are distributed uniformly on $[K]$. For $D_{\mathbf{Q}},D_{\mathbf{P}}\in\mathcal{D}$ with $\mathbf{Q}\neq\mathbf{P}$, we have that
@@
Consequently, we have that there exists a family of distributions of size $2^{\Omega(d^{c})}$ that is $(d^{\Omega(K)(c-1/2)}K,K)$ correlated. Hence, by C.3 we have that any SQ algorithm that solves the decision problem $\mathcal{B}(\mathcal{D},D)$ requires $2^{\Omega(d^{c})}d^{\Omega(K)(c-1/2)}=2^{d^{\Omega(1)}}$ (where we used the assumption that $d/\log(d)$ is much larger that $K$) queries to the $\text{VSTAT}(d^{\Omega(K)}$.
In order to complete the proof of the theorem, we need to show the following standard reduction from the testing problem to the learning problem.